OpenStudy (anonymous):

World oil consumption was 82.459 million barrels per day in 2005 and is increasing by about 1.3% per year. Let c_n be daily world oil consumption n years after 2005.

OpenStudy (fanduekisses):

Use the formula A=Pert Where P isthe principal (starting amount) r is rate and t is time

OpenStudy (anonymous):

I was able to figure that out I am unable to figure out what,, the \[\sum_{1}^{18}365c_n\] represents

OpenStudy (agreene):

that notation is giving you a sum: \[\sum_{i=1}^{n}i=\frac{n(n+1)}2\] in general terms

OpenStudy (agreene):

So, for this example: \[\sum_{C_n=1}^{18}365C_n=365\sum_{C_n}^{18}C_n\] \[365\left( \frac{18(19)}2 \right)\]